OpenStudy (anonymous):

Simplify completely quantity 8 x minus 16 over quantity x squared minus 13 x plus 22 and find the restrictions on the variable.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\huge\frac{8x-16}{x^2-13x+22}\]?

OpenStudy (anonymous):

I got x-5/x+1 as my answer but I don't know what the restriction would be

OpenStudy (misty1212):

did you factor the denomiantor?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry I meant 8/x-11

OpenStudy (misty1212):

\[x^2-13x+22=(x-2)(x-11)\]

OpenStudy (triciaal):

factor the denominator again hint 2*11 = 22 and 2 + 11 = 13

OpenStudy (misty1212):

then cancel and you get what you wrote

OpenStudy (anonymous):

But what would be the restrictions?

OpenStudy (misty1212):

so your answer is correct as for the "restrictions" the denominator cannot be zero

OpenStudy (misty1212):

\[(x-2)(x-11)=0\\ x=2,x=11\]

OpenStudy (misty1212):

so the restrictions are \(x\neq2\) and \(x\neq 11\) because if \(x\) where either of those two numbers, the denominator would be zero, and if you try to divide by zero you will explode

OpenStudy (anonymous):

Ok. I need to ask another one

OpenStudy (anonymous):

Simplify completely quantity x squared plus 4 x minus 45 all over x squared plus 10 x plus 9 and find the restrictions on the variable.

OpenStudy (misty1212):

lol the equation tool below is very useful

OpenStudy (anonymous):

I got x-5/x+1

OpenStudy (anonymous):

Sorry

OpenStudy (misty1212):

\[\huge \frac{x^2+4x-45}{x^2+10x+9}\]

OpenStudy (misty1212):

your answer is correct

OpenStudy (anonymous):

yes:) What would be the restrictions for this one? My answer is x-5/x+1

OpenStudy (misty1212):

repeat as before take the denominator in factored form set it equal to zero, solve for \(x\)

OpenStudy (anonymous):

x=-1?

OpenStudy (misty1212):

you obviously already factored the denominator as \[(x+1)(x+9)\] solve \[(x+1)(x+9)=0\] and solve for \(x\)

OpenStudy (anonymous):

-1,-9

OpenStudy (misty1212):

bingo

OpenStudy (misty1212):

that is, those are the restricted values

OpenStudy (anonymous):

ohh thank you:)

OpenStudy (anonymous):

That's all for now

OpenStudy (anonymous):

Simplify completely quantity 6 x minus 12 over 10.

OpenStudy (anonymous):

is it 3(x-5)/5